I hope that you can explain step by step, I am not very clear about some calculation formula, thank you!

Maggies Magazines (MM) has straight nonconvertible bond that currently yield 7%. MMs stock sells for $22 per share, has anWill call the issue in the first year that the conversion Conversion Value Straight debt Cash flow to value of bondholder ifConversion year = Value in conversion = b. What is the expected rate of return (i.e., before-tax component cost) on the propo
Answer #1
.(1)&(2) Anticipated stock Price & Conversion Value
Current stock Price=P0 $22
Growth Rateof stock price=7% 0.07
Stock Price in Year 1=22*(1+0.07) $23.54
Stock Price in Year (N+1)=1.07*(Stock Price in Year (N)
A B=A*32
Year Stock Price Conversion Value
0 $22 $704.00
1 $23.54 $753.28
2 $25.19 $806.01
3 $26.95 $862.43
4 $28.84 $922.80
5 $30.86 $987.40
6 $33.02 $1,056.51
7 $35.33 $1,130.47
8 $37.80 $1,209.60
9 $40.45 $1,294.28
10 $43.28 $1,384.87
11 $46.31 $1,481.82
12 $49.55 $1,585.54
13 $53.02 $1,696.53
14 $56.73 $1,815.29
15 $60.70 $1,942.36
16 $64.95 $2,078.32
17 $69.49 $2,223.81
18 $74.36 $2,379.47
19 $79.56 $2,546.04
20 $85.13 $2,724.26
Bonds par value $1,000
Conversion value at the time of convesrion $1,250.00 (1+0.25)*1000
YEAR OF CONVERSION 9
Conversion value at conversion $1,294.28
Share Value at conversion $40.45
.(3)and (4) STRAIGHT BOND PRICE and CASH Flow if Converted
Straight Bond Price =Present Value of future Cash Flow
Current Bond Price:
Pmt Coupon Payment per year $60.00 (6%*1000)
Nper Number of years of payment left 20
Rate Yield 7%
Fv Terminal Cash Flow $1,000
P0 Straight Bond Price in year 0 $894.06 (Using PV function of excel with Rate=7%,Nper=20, Pmt=-60, Fv=-1000)
Pn T(20-N) Number of years to maturity Straight Bond Price (Using PV function of excel with Rate=796, Nper=T, Pmt=-60, Fv=-10b Pv Nper Pmt Fv RATE EXPECTED RETURN WITH ANTICIPATED CONVERSION IN YEAR 9 Investment $894.06 (Note, the investment of invesH45 X fic A B C K L M N O P =PV(H43,442,-H41,-444) D E I J Conversion value at the time of convesrion $1,250.00 (1+0.25)*1000Clipboard Font Allghment Number Styles Lells Edith 152 x fc =PV(7%,H52,-60,-1000) E F A B C D I J K L M N O P Pmt Nper CurrenH76 * fic : IB =RATE(H73,474,-H72,H75) D E F C L M N O P HT 8 $940.29 P12 7 $946.11 P13 6 $952.33 P14 5 $959.00 P15 4 $966.13